OpenStudy (m_m13):

What is the value of x? will medal 1. x+4=-9 2. 12- x=7 3. -5x=-35 4. x/4-5=6 5. 3(x-2)+5x=18 6. 8x-9=3x=16 7. -2/7+6+1/7x+18

OpenStudy (chantysquirrel1129**):

One question per post please lol c:

OpenStudy (anonymous):

1. -14. this is because if you take -14 and add +4 then it equals -9. Try to say what in place of the x. #2. 12-what=7? I hope this helped

OpenStudy (texaschic101):

1. x + 4 = -9 --- subtract 4 from both sides x + 4 - 4 = -9 - 4 x = - 13 2. 12 - x = 7 --- subtract 12 from both sides 12 - 12 - x = 7 - 12 -x = - 5 -- multiply by -1 to make x positive -1(-x) = -1(-5) x = 5 3. -5x = -35 --- divide both sides by -5 (-5/-5)x = -35/-5 x = 7 4. x/4 - 5 = 6 --- add 5 to both sides x/4 - 5 + 5 = 6 + 5 x/4 = 11 --- multiply both sides by 4 4(x/4) = 4(11) x = 44 5. 3(x-2) + 5x = 18 -- distribute through the parenthesis 3x - 6 + 5x = 18 --- add 6 to both sides 3x - 6 + 6 + 5x = 18 + 6 3x + 5x = 18 + 6 8x = 24 -- divide both sides by 8 (8/8)x = 24/8 x = 3 6. you need to write this one over...there is a mistake in the problem 7. -2/7 + 6 + 1/7x + 18 --- combine like terms (-2/7 + 6 + 18) + 1/7x (-2/7 + 42/7 + 126/7) + 1/7x 166/7 + 1/7x....that's all you can do with this because there is no equal sign.